Cashis11:

Six machines can polish 42 tables in one hour. What is the ratio of the number of machines to the number of tables polished in one hour? 1:7 1:6 6:7 7:10

DanGonz:

6:7 bro

Cashis11:

thanks

DanGonz:

no rob can i get best response

Cashis11:

there you go

DanGonz:

thx

jhonyy9:

6 machines polish 42 tables what is the ratio ?

jhonyy9:

ratio mean 6/42 what simplified will be 1/7 yes ?

jhonyy9:

@supie any idea pls. ?

supie:

Yes, I agree \(\LARGE\frac{6}{42}\) would be incorrect because it has to be simplified \(\LARGE\frac{6}{42} \div \frac{6}{6}=\frac{1}{7}\) And a fraction is the same as a ratio.

sophia333:

@bella345

bella345:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sophia333 @bella345 \(\color{#0cbb34}{\text{End of Quote}}\) what

Jaylenn123:

Charlotte completed an experiment in which she investigated and discovered that an electrically charged object can attract an uncharged object without any contact. She used a PVC pipe and small pieces of paper for her experiment. Each time, she rubbed the PVC pipe with a cloth to give it a charge and then held it over paper. What would be Charlotte's control group in this experiment?